A car initially traveling at 43 m/s accelerates to 53 m/s. While speeding up, the car travels 200 m. Calculate the acceleration.


1
Expert's answer
2020-10-01T10:08:01-0400

The acceleration is


"a=\frac{v^2_\text{final}-v^2_\text{initial}}{2d}=\frac{53^2-43^2}{2\cdot200}=2.4\text{ m/s}^2."